OpenStudy (anonymous):

how can i solve this eq 3sech^2 x + 4tanh x = -1

OpenStudy (er.mohd.amir):

sec h^2 x=1-tan h^2x

OpenStudy (solomonzelman):

yes, use that, and let tanh(x)=a...

OpenStudy (anonymous):

i think i get the idea, let me try first, thank you

OpenStudy (solomonzelman):

And then, you can use these two rules \(\color{#000000 }{ \displaystyle \sinh(x)=\frac{e^x-e^{-x}}{2} }\) \(\color{#000000 }{ \displaystyle \cosh(x)=\frac{e^x+e^{-x}}{2} }\) to solve for x.

OpenStudy (solomonzelman):

(You can extract other hyperbolic functions from these)

OpenStudy (anonymous):

i got the answer correctly. thank you so much to both of you

OpenStudy (solomonzelman):

You Welcome!